# Enumerating Finite Braid Group Orbits on $SL_{2}(\mathbb{C})$-Character Varieties

### v2

@@ 5.1.4 $G_{26}$
The group $G_{26}$ of order 1296 is also denoted $W(M_{3})$ with degrees $\deg(G_{26})=\{6,12,18\}$ [21, Table D.3]. In $G_{26}$ there are three conjugacy classes of reflections, corresponding to nontrivial eigenvalues in $\{\zeta_{3},\zeta_{3}^{2},-1\}$. In this group the are 6 sets of eigenvalues for the inverse product, which correspond to 3 inverse pairs, so we only list one from each pair.
@@ A $(w,w^{2},-1)$$\{\frac{1}{12},\frac{7}{12},\frac{5}{6}\}$
B $(w,w,-1)$$\left\{\frac{5}{18},\frac{11}{18},\frac{17}{18}\right\}$
@@ C $(w^{2},w^{2},-1)$$\{\frac{1}{6},\frac{1}{6},\frac{5}{6}\}$
In $G_{26}$ choice of different exemplars nor the inverses for any type produce new orbits. The middle convolution data for this group can be found in Table 28. Further, all three orbits from Type A are distinct, as are all three orbits from Type B. When comparing with Tykhyy’s classification, there was only one orbit (from Type B) that we could not find an exact match for. We listed the closest match, which seemed to be Orbit 5 in [31] and distinguish the differing coordinate with an asterisk.
@@
$\lambda$ $\omega_{X,Y,Z,4}$ Lisovy-Tykhyy ##\## $\theta_{1,2,3,4}$ $\sigma_{12,23,13,24}$
Tykhyy ##\##
@@
$-z^{3}$ $(0,-\sqrt{3},0,0)$ Type II, $X^{\prime}=0,X^{\prime\prime}=-\sqrt{3}$ $(\frac{1}{4},\frac{1}{12},\frac{1}{4},\frac{11}{12})$ $(\frac{1}{2},\frac{1}{2},\frac{1}{2},\frac{5}{6})$ 4 2 24 0
$-z^{6}+1$ $(0,-z^{9}+2z^{3},0,0)$ Type II, $X^{\prime}=0,X^{\prime\prime}=-z^{9}+2z^{3}$ $(\frac{1}{6},\frac{1}{2},\frac{2}{3},\frac{1}{2})$ $(\frac{1}{2},\frac{1}{2},,\frac{5}{6},\frac{1}{2})$ 3 2 24 $\langle 24,4\rangle$
@@
$-z^{4}$ $(z^{9}-2z^{3},z^{9}-2z^{3},2,-1)$ Type III, $\omega=z^{9}-2z^{3}$ $(\frac{1}{9},\frac{1}{18},\frac{1}{18},\frac{2}{3})$ $(\frac{1}{2},\frac{1}{3},\frac{1}{2},\frac{1}{6})$ 5 3 36 0
$-z^{16}$ $(-z^{9}+2z^{3},-z^{9}+2z^{3},2,-1)$ Type III, $\omega=-z^{9}+2z^{3}$ $(\frac{5}{9},\frac{5}{18},\frac{5}{18},\frac{2}{3})$ $(\frac{1}{2},\frac{1}{3},\frac{1}{2},\frac{1}{6})$ 5 3 36 0
@@
$-z^{6}+1$ $(0,0,2,-1)$ Type III, $\omega=0$ $(1,\frac{1}{2},\frac{1}{2},\frac{2}{3})$ $(\frac{1}{2},\frac{1}{3},\frac{1}{2},\frac{1}{2})$ 5 3 12 $\langle 12,1\rangle$
Table 8: Here $w=\zeta_{3},z=\zeta_{36}$. The correct signature for Tykhyy’s orbit $5$ would have $3(\frac{7}{18})\equiv\frac{1}{6}$ as $\sigma_{24}$ but such a signature was not in the orbit so we provide the closest match. The subgroups $\langle 24,4\rangle$ and $\langle 12,1\rangle$ are the dicyclic groups of order 24 and 12 respectively.
@@ 5.1.5 $G_{27}$
A $(-1,-1-1)$ $\{\frac{2}{60},\frac{38}{60},\frac{50}{60}\}$
B $(-1,-1-1)$ $\left\{\frac{5}{60},\frac{35}{60},\frac{50}{60}\right\}$
C $(-1,-1,-1)$ $\{\frac{10}{60},\frac{34}{60},\frac{46}{60}\}$
@@
In $G_{27}$ for all three types, different exemplars do not produce new orbits. Moreover the inverse types do not produce new orbits and we have checked that these 9 orbits are distinct. The middle convolution details for Type A, B and C are listed in Table 31, 31, and 31 respectively. The comparison data is provided in Table 10. In this group we also find nice 4-tuples, details provided in Section 5.2.4
@@
$-z^{8}$ $(\frac{3-\sqrt{5}}{2},\frac{3-\sqrt{5}}{2},\frac{3-\sqrt{5}}{2},\sqrt{5}-1)$ 26 $(\frac{13}{15},\frac{1}{5},\frac{2}{15},\frac{2}{15})$ $(\frac{3}{5},\frac{2}{5},\frac{1}{2},\frac{1}{2})$ 87 15 60 0
$-z^{20}$ $(\frac{3-\sqrt{5}}{2},\frac{3-\sqrt{5}}{2},\frac{3-\sqrt{5}}{2},\sqrt{5}-1)$ 26 $(\frac{1}{3},\frac{1}{3},\frac{3}{5},\frac{1}{3})$ $(\frac{3}{5},\frac{3}{5},\frac{1}{2},\frac{1}{2})$ 81 15 60 $\langle 120,5\rangle$
@@
$-z^{5}$ (1, 1, 1, 1) 39 $(\frac{3}{4},\frac{1}{12},\frac{1}{12},\frac{1}{12})$ $(\frac{1}{2},\frac{3}{5},\frac{1}{3},\frac{1}{2})$ 114 24 96 0
$-z^{20}$ (1, 1, 1, 1) 39 $(\frac{1}{2},\frac{1}{3},\frac{2}{3},\frac{1}{3})$ $(\frac{1}{2},\frac{2}{5},\frac{1}{2},\frac{1}{3})$ 112 24 96 $\langle 120,5\rangle$
@@
$-z^{4}$ $(\frac{3+\sqrt{5}}{2},\frac{3+\sqrt{5}}{2},\frac{3+\sqrt{5}}{2},-1-\sqrt{5})$ 27 $(\frac{14}{15},\frac{14}{15},\frac{3}{5},\frac{1}{15})$ $(\frac{1}{5},\frac{4}{5},\frac{1}{2},\frac{1}{2})$ 83 15 60 0
$-z^{16}$ $(\frac{3+\sqrt{5}}{2},\frac{3+\sqrt{5}}{2},\frac{3+\sqrt{5}}{2},-1-\sqrt{5})$ 27 $(\frac{4}{15},\frac{4}{15},\frac{2}{5},\frac{4}{15})$ $(\frac{1}{5},\frac{1}{5},\frac{1}{2},\frac{1}{2})$ 82 15 60 0
@@ 5.2 Nice 4-Tuples in Rank 3 Groups
In comparing with the existing classifications, we can only compare with Tykhyy’s $\mathcal{M}^{(5)}$ orbits [31] since our 4-tuples gives us local systems on the 5-punctured sphere. Unfortunately, without the $\omega$ parameters from the work of Lisovy-Tykhyy [22] like in the 3-tuple cases, it is more difficult to relate the our orbit sizes to Tykhyy’s. We still computed all the signatures for the different possible equivalent representatives for each orbit and attempted to find a match, which was successful.
@@ A $(-1,-1,-1,-1)$ $\{\frac{1}{2},\frac{1}{2},0\}$
The middle convolution details can be found in Table B. This type is its own inverse. For this tuple we could not compute the orbit size as it was too large, but we were able to find a signature in the orbit (up to Tykhyy’s equivalence relations) that agreed with one listed in Tykhyy’s paper. Because we cannot compute the full braid group orbit, we also cannot check if different exemplars would result in a different orbit after middle convolution. We provide the comparison details in Table 12.
@@
A $-1$ $(1,1,1,1,1)$ $(\frac{1}{3},\frac{1}{3},\frac{1}{3},\frac{1}{3},\frac{1}{3},\frac{1}{5},\frac{1}{5})$ 8 192 - 0
@@ C $(z^{2},z,z,z^{2})$$\{\frac{2}{3},\frac{1}{6},\frac{1}{6}\}$
The middle convolution details for $G_{25}$ can be found in Table B. The orbits for Type A and Type B are the same. All three types do not split into new orbits and their inverses do not produce new orbits. The comparison details for the two distinct orbits can be found in Table 14.
@@ 5.2.4 $G_{27}$
A $(-1,-1,-1,-1)$ $\{\frac{5}{6},\frac{5}{6},\frac{1}{3}\}$
@@
A $-z^{5}$ $(\frac{1}{3},\frac{1}{3},\frac{1}{3},\frac{1}{3},\frac{1}{3})$ $(0,\frac{1}{5},\frac{1}{3},\frac{1}{3},\frac{1}{5},\frac{3}{5},\frac{1}{3})$ 87 432 - $\langle 120,5\rangle$
@@ 5.3 Nice 4-Tuples in Rank 4 Groups
In this case we are looking for nice tuples of 4 reflections generating the whole group such that their inverse product has a nontrivial eigenvalue of $T-2=2$. Of the five rank 4 reflection groups, we found nice 4-tuples in $G_{28}$, $G_{30}$ and $G_{32}$. In the case of $G_{32}$, the field containing all the eigenvalues would be $\mathbb{Q}(\zeta_{360})$ which is much too large to perform the computations over. For $G_{32}$, in Section 5.3.3 we justify why it is enough to search for such tuples over $\mathbb{Q}(\zeta_{12})$.
After taking the middle convolution, multiplying with a suitable character and adjoining the inverse product, our tuples become $SL_{2}(\mathbb{C})$ local systems on the 5-punctured sphere. We found that in some cases, one of the 5 matrices in the local system ends up being the identity, which means our tuple comes from some local system on the 4-punctured sphere. In the case where we have 4 nontrivial matrices, we can compute the $\omega$ parameters in the notation of Lisovy-Tykhyy [22] and identify the corresponding $\mathcal{M}^{(4)}$ orbit in Tykhyy’s paper [31]. When all 5 of the matrices are nontrivial, we can only compare with Tykhyy’s $\mathcal{M}^{(5)}$ orbit signatures from [31].
@@ 5.3.1 $G_{28}$
The group $G_{28}$ of order 1152 is also denoted $W(F_{4})$ with degrees $\deg(G_{28})=\{2,6,8,12\}$ [21, Table D.3]. All reflections in $G_{28}$ have nontrivial eigenvalue -1, and form two conjugacy classes of reflections. We only found one type of nice 4-tuples in this group, with two possible choices of $\lambda$ parameters for middle convolution.
@@
A $(-1,-1,-1,-1)$ $\{\frac{1}{6},\frac{1}{6},\frac{5}{6},\frac{5}{6}\}$
@@
A different choice of exemplar does not produce a new orbit and the two orbits computed above are the same, so we only get one new orbit from this group. The middle convolution details can be found in Table 34. After putting the matrices in $SL_{2}(\mathbb{C})$, we see that the inverse product of the middle convolution of this tuple is the identity, so this tuple comes from a local system on the 4-punctured sphere. As in the case of 3-tuples, we provide the details comparing with the results of both Lisovy-Tykhyy and Tykhyy ([22], [31]) in Table 20. The $\omega$ parameters appear to correspond to Orbit 31, but in fact this orbit is of Type III, as described in Lemma 39 in [22].
@@
$z^{4}$ $(2,2,2,-1)$ Type III, $\omega=2$ $(\frac{2}{3},\frac{1}{3},\frac{1}{3},\frac{2}{3})$ $(\frac{1}{2},\frac{1}{3},\frac{1}{2},1)$ 5 3 9 $\langle 24,3\rangle$

### v3

@@ 5.1.4 $G_{26}$
The group $G_{26}$ of order 1296 is also denoted $W(M_{3})$ with degrees $\deg(G_{26})=\{6,12,18\}$ [ld2009, Table D.3]. In $G_{26}$ there are three conjugacy classes of reflections, corresponding to nontrivial eigenvalues in $\{\zeta_{3},\zeta_{3}^{2},-1\}$. In this group the are 6 sets of eigenvalues for the inverse product, which correspond to 3 inverse pairs, so we only list one from each pair.
@@ A $(w,w^{2},-1)$ $\{\frac{1}{12},\frac{7}{12},\frac{5}{6}\}$
B (w,w,1)(w,w,-1) $\left\{\frac{5}{18},\frac{11}{18},\frac{17}{18}\right\}$
@@ C $(w^{2},w^{2},-1)$ $\{\frac{1}{6},\frac{1}{6},\frac{5}{6}\}$
In $G_{26}$ choice of different exemplars nor the inverses for any type produce new orbits. The middle convolution data for this group can be found in Table 34. Further, all three orbits from Type A are distinct, as are all three orbits from Type B. When comparing with Tykhyy’s classification, there were two orbit (from Type B) that we could not find an exact match for. We listed the closest match, distinguishing the differing coordinate with an asterisk. Note that both cases agree with Orbit 5 in [Tyk, Table 1] when the formulas are read modulo 2 and up to sign, as Tykhyy intended.
@@
$\lambda$ $\omega_{X,Y,Z,4}$ Lisovyy-Tykhyy #\# $\theta_{1,2,3,4}$ $\sigma_{12,23,13,24}$
Tykhyy #\#
@@
$z^{3}$ $(0,-\sqrt{3},0,0)$ Type II, $X^{\prime}=0,X^{\prime\prime}=-\sqrt{3}$ $(\frac{1}{4},\frac{1}{12},\frac{1}{4},\frac{11}{12})$ $(\frac{1}{2},\frac{1}{2},\frac{1}{2},\frac{5}{6})$ 4 2 24 0
$z^{6}+1$ $(0,-z^{9}+2z^{3},0,0)$ Type II, $X^{\prime}=0,X^{\prime\prime}=-z^{9}+2z^{3}$ $(\frac{1}{6},\frac{1}{2},\frac{2}{3},\frac{1}{2})$ $(\frac{1}{2},\frac{1}{2},,\frac{5}{6},\frac{1}{2})$ 3 2 24 $\langle 24,4\rangle$
@@
$z^{4}$ $(z^{9}-2z^{3},z^{9}-2z^{3},2,-1)$ Type III, $\omega=z^{9}-2z^{3}$ $(\frac{1}{9},\frac{1}{18},\frac{1}{18},\frac{2}{3})$ $(\frac{1}{2},\frac{1}{3},\frac{1}{2},\frac{1}{6})$ 5 3 36 0
$z^{16}$ $(-z^{9}+2z^{3},-z^{9}+2z^{3},2,-1)$ Type III, $\omega=-z^{9}+2z^{3}$ $(\frac{5}{9},\frac{5}{18},\frac{5}{18},\frac{2}{3})$ $(\frac{1}{2},\frac{1}{3},\frac{1}{2},\frac{1}{6}^{*})$ $5^{*}$ 3 36 0
@@
$z^{6}+1$ (0,0,2,1)(0,0,2,-1) Type III, $\omega=0$ $(1,\frac{1}{2},\frac{1}{2},\frac{2}{3})$ $(\frac{1}{2},\frac{1}{3},\frac{1}{2},\frac{1}{2})$ 5 3 12 $\langle 12,1\rangle$
Table 12: Here $w=\zeta_{3},z=\zeta_{36}$. The subgroups $\langle 24,4\rangle$ and $\langle 12,1\rangle$ are the dicyclic groups of order 24 and 12 respectively. The signature for Tykhyy’s orbit $5$ is $\theta_{1,2,3,4}=(2x,x,x,\frac{2}{3})$ and $\sigma_{12,23,13,24}=(\frac{1}{2},\frac{1}{3},\frac{1}{2},3x)$ but such a signature was not in the orbit so we provide the closest match. The coordinate that differs from the formula is marked with an asterisk, but note that as $\sigma_{24}$ is only defined modulo 2 and up to sign, we believe this still corresponds to Orbit 5 of Tykhyy’s Table 1.
@@ 5.1.5 $G_{27}$
A (1,11)(-1,-1-1) $\{\frac{2}{60},\frac{38}{60},\frac{50}{60}\}$
B (1,11)(-1,-1-1) $\left\{\frac{5}{60},\frac{35}{60},\frac{50}{60}\right\}$
C (1,1,1)(-1,-1,-1) $\{\frac{10}{60},\frac{34}{60},\frac{46}{60}\}$
@@
In $G_{27}$ for all three types, different exemplars do not produce new orbits. Moreover the inverse types do not produce new orbits and we have checked that these 9 orbits are distinct. The middle convolution details for Type A, B and C are listed in Table 37, 37, and 37 respectively. The comparison data is provided in Table 14. In this group we also find nice 4-tuples, details provided in Section 5.2.4
@@
$z^{8}$ $(\frac{3-\sqrt{5}}{2},\frac{3-\sqrt{5}}{2},\frac{3-\sqrt{5}}{2},\sqrt{5}-1)$ 26 $(\frac{13}{15},\frac{1}{5},\frac{2}{15},\frac{2}{15})$ $(\frac{3}{5},\frac{2}{5},\frac{1}{2},\frac{1}{2})$ 87 15 60 0
$z^{20}$ $(\frac{3-\sqrt{5}}{2},\frac{3-\sqrt{5}}{2},\frac{3-\sqrt{5}}{2},\sqrt{5}-1)$ 26 $(\frac{1}{3},\frac{1}{3},\frac{3}{5},\frac{1}{3})$ $(\frac{3}{5},\frac{3}{5},\frac{1}{2},\frac{1}{2})$ 81 15 60 $\langle 120,5\rangle$
@@
$z^{5}$ (1, 1, 1, 1) 39 $(\frac{3}{4},\frac{1}{12},\frac{1}{12},\frac{1}{12})$ $(\frac{1}{2},\frac{3}{5},\frac{1}{3},\frac{1}{2})$ 114 24 96 0
$z^{20}$ (1, 1, 1, 1) 39 $(\frac{1}{2},\frac{1}{3},\frac{2}{3},\frac{1}{3})$ $(\frac{1}{2},\frac{2}{5},\frac{1}{2},\frac{1}{3})$ 112 24 96 $\langle 120,5\rangle$
@@
$z^{4}$ $(\frac{3+\sqrt{5}}{2},\frac{3+\sqrt{5}}{2},\frac{3+\sqrt{5}}{2},-1-\sqrt{5})$ 27 $(\frac{14}{15},\frac{14}{15},\frac{3}{5},\frac{1}{15})$ $(\frac{1}{5},\frac{4}{5},\frac{1}{2},\frac{1}{2})$ 83 15 60 0
$z^{16}$ $(\frac{3+\sqrt{5}}{2},\frac{3+\sqrt{5}}{2},\frac{3+\sqrt{5}}{2},-1-\sqrt{5})$ 27 $(\frac{4}{15},\frac{4}{15},\frac{2}{5},\frac{4}{15})$ $(\frac{1}{5},\frac{1}{5},\frac{1}{2},\frac{1}{2})$ 82 15 60 0
@@ 5.2 Nice 4-Tuples in Rank 3 Groups
In comparing with the existing classifications, we can only compare with Tykhyy’s $\mathcal{M}^{(5)}$ orbits [Tyk] since our 4-tuples gives us local systems on the 5-punctured sphere. Unfortunately, without the $\omega$ parameters from the work of Lisovyy-Tykhyy [T-L] like in the 3-tuple cases, it is more difficult to relate our orbit sizes to Tykhyy’s. We still computed all the signatures for the different possible equivalent representatives for each orbit and attempted to find a match, which was successful.
@@ 5.2.1 $G_{23}$
The middle convolution details can be found in Table 38. This type is its own inverse. For this tuple we could not compute the orbit size as it was too large, but we were able to find a signature in the orbit (up to Tykhyy’s equivalence relations) that agreed with one listed in Tykhyy’s paper. Because we cannot compute the full braid group orbit, we also cannot check if different exemplars would result in a different orbit after middle convolution. We provide the comparison details in Table 16.
@@
A 1-1 $(1,1,1,1,1)$ $(\frac{1}{3},\frac{1}{3},\frac{1}{3},\frac{1}{3},\frac{1}{3},\frac{1}{5},\frac{1}{5})$ 8 192 - 0
@@ C $(z^{2},z,z,z^{2})$ $\{\frac{2}{3},\frac{1}{6},\frac{1}{6}\}$
The middle convolution details for $G_{25}$ can be found in Table 39. The orbits for Type A and Type B are the same. All three types do not split into new orbits and their inverses do not produce new orbits. The comparison details for the two distinct orbits can be found in Table 18.
@@ 5.2.4 $G_{27}$
A (1,1,1,1)(-1,-1,-1,-1) $\{\frac{5}{6},\frac{5}{6},\frac{1}{3}\}$
@@
A $z^{5}$ $(\frac{1}{3},\frac{1}{3},\frac{1}{3},\frac{1}{3},\frac{1}{3})$ $(0,\frac{1}{5},\frac{1}{3},\frac{1}{3},\frac{1}{5},\frac{3}{5},\frac{1}{3})$ 87 432 - $\langle 120,5\rangle$
@@ 5.3 Nice 4-Tuples in Rank 4 Groups
In this case we are looking for nice tuples of 4 reflections generating the whole group such that their inverse product has a nontrivial eigenvalue of T2=2T-2=2. Of the five rank 4 reflection groups, we found nice 4-tuples in $G_{28}$, $G_{30}$ and $G_{32}$. In the case of $G_{32}$, the field containing all the eigenvalues would be $\mathbb{Q}(\zeta_{360})$ which is much too large to perform the computations over. For $G_{32}$, in Section 5.3.3 we justify why it is enough to search for such tuples over $\mathbb{Q}(\zeta_{12})$.
After taking the middle convolution, multiplying with a suitable character and adjoining the inverse product, our tuples become $SL_{2}(\mathbb{C})$ local systems on the 5-punctured sphere. We found that in some cases, one of the 5 matrices in the local system ends up being the identity, which means our tuple comes from some local system on the 4-punctured sphere. In the case where we have 4 nontrivial matrices, we can compute the $\omega$ parameters in the notation of Lisovyy-Tykhyy [T-L] and identify the corresponding $\mathcal{M}^{(4)}$ orbit in Tykhyy’s paper [Tyk]. When all 5 of the matrices are nontrivial, we can only compare with Tykhyy’s $\mathcal{M}^{(5)}$ orbit signatures from [Tyk].
@@ 5.3.1 $G_{28}$
The group $G_{28}$ of order 1152 is also denoted $W(F_{4})$ with degrees $\deg(G_{28})=\{2,6,8,12\}$ [ld2009, Table D.3]. All reflections in $G_{28}$ have nontrivial eigenvalue -1, and form two conjugacy classes of reflections. We only found one type of nice 4-tuples in this group, with two possible choices of $\lambda$ parameters for middle convolution.
@@
A (1,1,1,1)(-1,-1,-1,-1) $\{\frac{1}{6},\frac{1}{6},\frac{5}{6},\frac{5}{6}\}$
@@
A different choice of exemplar does not produce a new orbit and the two orbits computed above are the same, so we only get one new orbit from this group. The middle convolution details can be found in Table 42. After putting the matrices in $SL_{2}(\mathbb{C})$, we see that the inverse product of the middle convolution of this tuple is the identity, so this tuple comes from a local system on the 4-punctured sphere. As in the case of 3-tuples, we provide the details comparing with the results of both Lisovyy-Tykhyy and Tykhyy ([T-L], [Tyk]) in Table 24. The $\omega$ parameters appear to correspond to Orbit 31, but in fact this orbit is of Type III, as described in Lemma 39 in [T-L].
@@
$z^{4}$ (2,2,2,1)(2,2,2,-1) Type III, $\omega=2$ $(\frac{2}{3},\frac{1}{3},\frac{1}{3},\frac{2}{3})$ $(\frac{1}{2},\frac{1}{3},\frac{1}{2},1)$ 5 3 9 $\langle 24,3\rangle$